Problem 6. You have a uniform string with a mass of 0.0130 kg and length 1.75 m under a tension 10.0 N. The string is fixed at both ends, and is vibrating at its fourth resonant frequency (i.e. the fourth harmonic). a) (2 points) What is the wavelength of the standing wave in the string? b) (2 points) What is the frequency?
Answer #1

Velocity ven 0 Co 0135/1.ts) m /L、 = 367 mls 2 L 2x).75 1.93再 CA,s) 41.93
Answer #2

Wavelength

For a standing transverse wave in a stretched string that is attached at both ends, the number of the normal mode equals the number of half wavelengths that are contained in its length. The wavelength λn of the nth normal mode is therefore

λn=2Lnn=1,2,3,...

where L denotes the string's length.

In the question, you are given L and n. Substitute to obtain the numerical answer for the wavelength.

λ4=2(1.91 m)4=0.955 m

Frequency

The frequency of the standing wave in the nth normal mode, which is the nth harmonic fn, can be found using the relation that is valid for all periodic waves

fn=vλn

where λn is the wavelength of the nth normal mode (see the formula for wavelength) and v is the speed of transverse waves in the string. This speed is not given directly, but can be obtained from the formula

v=Fm/L

where F is the given tension in the string and m and L are its given mass and length, respectively. Perform all necessary substitutions to obtain the frequency f4 in terms of given quantities.

fn=n2LFm/Ln=1,2,3,...

For the value of the fourth harmonic,




answered by: Muhammad Aslam